OpenStudy (anonymous):

Which rule will find the nth term of an arithmetic sequence with a rate of increase equal to 1/2 and a6 = 18? a.) an = 30 + n/2 b.) an= 7+ n/2 c.) an= 15- n/n d.) an= 15+ n/2 Thank you for your help! ^_^ xoxo

OpenStudy (zehanz):

Possible first step: set n=6 in these formulas to filter out the wrong ones.

rishavraj (rishavraj):

\[a_n = a_1 + (n - 1)d\] u r having a_6 find a_1 and plug it in the above equation...... and also the way @ZeHanz said would be far easy....

OpenStudy (anonymous):

how do I put the an on mathway.com? Lol I don't know how! ;-;

OpenStudy (zehanz):

Not sure what you mean... If I set n=6 in the first one, I get: 30 + 6/2 = 30+3=33, so wrong. Only if two of the formulas give you 18 as an aswer, you have to do extra work (or thinking...)